A boy is 12 years older than his sister. If in 4 years time he is twice her age, what are their current ages?
Answer by Cromlix(4381)

Hi there,
A boy is 12 years older than his sister
B = 12 + S.........(10
If in 4 years time he is twice her age
B + 4 = 2(S + 4)
Replace B in Eq(2) with (12 + S) from Eq(1)
12+ S + 4 = 2S + 8
Collect like terms
2S - S = 12 + 4 - 8
S = 8
Sister is 8 years old
Brother is 20 years old
Hope this helps :-)